GlueX TOF Meeting, October 24, 2018
GlueX Time-of-Flight Meeting
Wednesday, October 24, 2018
10:00 am EDT
JLab: CEBAF Center, Room F226
BlueJeans Meeting number is 350 531 998
- Review of minutes from the October 10 meeting
- Two current issues with TOF
- Calibration Status
- Lucite Shield Geometry
- Upgrade Plan
- NIM Paper (Paul)
- Action Item Recap
Talks can be deposited in the directory
/group/halld/www/halldweb/html/talks/2018 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2018/ .
- JLab: Thomas Britton, Ashley Ernst, Mark Ito (chair), Sasha Ostrovidov, Simon Taylor, Beni Zihlmann
There is a recording of this meeting on the BlueJeans site. Use your JLab credentials to access it.
The JAG Tailgate Party is today from 3 pm to 6 pm on the field behind the Residence Facility.
Review of minutes from the October 10 meeting
We went over the minutes.
Lucite Shield Simulation
Thomas has transferred his knowledge on the HDDS geometry of the Lucite shield to Ashley. The main missing piece is the material definition. She has been busy with the calibration but will start on this soon.
Particle ID Working Group
Thomas has agreed to convene a meeting of the new Particle Identification Working group. He will make an announcement next week.
Two current issues with TOF
Beni and Sasha reported on a couple of areas of concern.
- TOF ADC config file. Beni reported several channels with FADC configuration/performance issues. See the log entry for details.
- Additional info. Sasha reports two problems: (1) intermittent drop out of FADC scaler channels and (2) strange baseline noise, intermittently on one of the channels Beni flagged in his log message. See his addendum to the Beni's log entry for details. Beni and Ashley will look at the channel on a scope tomorrow during the down time.
Ashley showed slides of the time difference fits from a recent run. She noted that there are a significant number of channels where the fit does not converge. Only 542 of 756 resulted in a peak within the fit range. She analyzed a run with four TOF skim files; Beni remarked that she needs more events; he recommends using a run with between 80 and 100 files. File:Tof 2018-08 fits run50638 part1.pdf
Upgrade Plan: Option #3
Sasha reported that Paul has been thinking about another option for re-configuring the inner region of the TOF. This involves having per plane 4 half-length counters (those that do not span the beam hole) of 3/4 width (4.5 cm wide) and 6 half-width counters (3.0 cm wide). Outside of those. See figures below (two versions of same design, constrasted with current design, shown crudely in both cases). This is option #3. It has 4 more PMTs per plane than the current array. Recall that option #2 has the same number of PMTs as the current array. This option does have a problem with accommodating the light guides for the three consecutive half-width counters; Paul is working on this. Sasha is planning to estimate rates for this configuration as well.